OpenStudy (anonymous):

quadratic equation (3x+2)(x^2-x-30)=0

hero (hero):

(3x+2)(x^2-6x+5x-30) =0 (3x+2)(x(x-6)+5(x-6))=0 (3x+2)(x-6)(x+5)=0 x = -5,-2/3, 6

hero (hero):

Any questions?

hero (hero):

If so, just post it here and I'll get back to you immediately.

OpenStudy (anonymous):

thanx i understand